Scores of an IQ test have a​ bell-shaped distribution with a mean of 100 and a standard deviation of 19. Use the empirical rule to determine the following. ​

(a) What percentage of people has an IQ score between 62 and 138​? ​

(b) What percentage of people has an IQ score less than 62 or greater than 138​? ​

(c) What percentage of people has an IQ score greater than 119​?
Answer #1

Given,

\mu = 100 , \sigma = 19

a)

z-score = (X - \mu ) / \sigma

For x = 62

z = ( 62 - 100) / 19

= -2

62 is 2 standard deviation below the mean.

For x = 138

z = ( 138 - 100) / 19

= 2

138 is 2 standard deviation above the mean.

Using empirical ( 68 - 95 - 99.7) rule,

Approximately 95% of the data falls in 2 standard deviation of the mean

P(62 < X < 138) = 95%

b)

P(X < 62 OR X > 138) = 1 - P(62 < X < 138)

= 1 - 0.95

= 0.05

= 5%

c)

z = ( 119 - 100) / 19

= 1

119 is 1 standard deviation above the mean.

Using empirical ( 68 - 95 - 99.7) rule,

Approximately 68% of the data falls in 1 standard deviation of the mean

That is ( 1 - 0.68) = 0.32 of the data falls outside the 1 standard deviation.

This means 16% data falls to either side of outside region of the 1 standard deviation.

P(X > 119) = 16%